### v2

# Dynamic Target Attack

Abstract
Existing gradient-based jailbreak attacks typically optimize an adversarial suffix to induce a fixed affirmative response, e.g.,Sure, here is…”. However, this fixed target usually resides in an extremely low-density region of a safety-aligned LLM’s output distribution conditioned on diverse harmful inputs. Due to the substantial discrepancy between the target and the original output, existing attacks require numerous iterations to optimize the adversarial prompt, which might still fail to induce the low-probability target response from the target LLM. In this paper, we propose Dynamic Target Attack (DTA), a new jailbreaking framework relying on the target LLM’s own responses as targets to optimize the adversarial prompts. In each optimization round, DTA iteratively samples multiple candidate responses directly from the output distribution conditioned on the current prompt, and selects the most harmful response as a temporary target for prompt optimization. In contrast to existing attacks, DTA significantly reduces the discrepancy between the target and the output distribution, substantially easing the optimization process to search for an effective adversarial prompt.
Extensive experiments demonstrate the superior effectiveness and efficiency of DTA: under the white-box setting, DTA only needs $200$ optimization iterations to achieve an average attack success rate (ASR) of over $87\%$ on recent safety-aligned LLMs, exceeding the state-of-the-art baselines by over $15\%$. The time cost of DTA is $\thicksim2$6 times less than existing baselines. Under the black-box setting, DTA uses Llama-3-8B-Instruct as a surrogate model for target sampling and achieves an ASR of $85\%$ against the black-box target model Llama-3-70B-Instruct, exceeding its counterparts by over $25\%$. All code and other materials are available here.
Warning: This paper may contain harmful or offensive content.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide spectrum of natural language tasks [1, 2, 3, 4]. To steer these capabilities toward beneficial and safe outputs, alignment techniques such as Reinforcement Learning from Human Feedback (RLHF) are employed to instill models with human values [5], which will significantly lower the probability of generating affirmative responses on harmful queries.
Despite the effectiveness of these alignment techniques, most existing gradient-based jailbreak methods [6, 7, 8] could jailbreak safety-aligned LLMs with a common paradigm, i.e., optimizing an adversarial suffix $S$ appended to a harmful prompt $P$ to maximize the probability of the target LLM generating a fixed affirmative response $T_{\text{fixed}}$. However, due to the alignment techniques, $T_{\text{fixed}}$ typically lies in an extremely low-density region of the LLM’s output distribution conditioned on $P+S$ (e.g., $\log p(T_{\text{fixed}}\mid P+S)=-52.63$ in Figure 1), creating a substantial discrepancy between the optimization objective (i.e., $T_{\text{fixed}}$) and the target LLM’s original outputs. As a result, these methods usually require thousands of iterations for optimizing $S^{*}$, while the probability of producing $T_{\text{fixed}}$ conditioned on $P+S^{*}$ remains low (e.g., $\log p(T_{\text{fixed}}\mid P+S^{*})=-39.07$ in Figure 1).
Figure 1: DTA directly samples $T_{\text{sampled}}$ from the LLM, which is more probable than $T_{\text{fixed}}$.
Our approach. We propose Dynamic Target Attack (DTA), a new jailbreaking framework that reframes target selection as an adaptive search over relatively high-density regions of the target LLM’s output space conditioned on the prompt, significantly reducing the discrepancy between the targets and LLM outputs. As shown in Figure 2, DTA (i) directly samples candidate harmful responses from high-density regions of the target LLM’s conditional distribution, (ii) selects the most harmful candidate as a temporary target ($T_{\text{sampled}}$) and performs a few optimization steps, and (iii) re-samples the target from the more harmful conditional distribution after optimization, iterating this sampling–optimization cycle. Due to the reduced discrepancy, the proposed sampling–optimization cycle substantially speeds the optimization process to identify an effective adversarial suffix. For example, in Figure 1, before optimization, the log-probability of generating the sampled response $T_{\text{sampled}}$ conditioned on $P+S$ ($\log p(T_{\text{sampled}}\mid P+S^{*})=-36.68$) is 1.4x higher than that of the fixed target $T_{\text{fixed}}$ ($\log p(T_{\text{fixed}}\mid P+S^{*})$=-52.63). After a few optimization steps, the conditional log-probability improves to 28.03-28.03. To further accelerate convergence, DTA then re-samples from the target LLM conditioned on the harmful prompt and the current updated suffix, which increases the chance of selecting an even more harmful, higher-density target for the next round and thereby further reduces the discrepancy between the target and the model’s output distribution.
We demonstrate that DTA effectively jailbreaks multiple recent safety-aligned LLMs, whether under the white-box setting or the black-box setting. Specifically, under the white-box setting, DTA only needs 200 optimization steps to achieve an average Attack Success Rate (ASR) of over 87% on five recent safety-aligned LLMs, which is over 15% higher than its counterparts. Furthermore, under the black-box setting, DTA effectively achieves an ASR of 85% against Llama-3-70B-Instruct, using Llama-3-8B-Instruct as the surrogate model, achieving an ASR of 85%. Our primary contributions can be summarized as follows:
  • We introduce Dynamic Target Attack (DTA), a novel jailbreak framework that effectively jailbreaks LLMs under both white- and black-box settings. In contrast to existing methods, DTA iteratively samples multiple candidate responses directly from the target model’s output distribution, and selects the most harmful as a temporary target for the current optimization round.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of DTA. For example, under the white-box setting, DTA requires only 200 optimization steps to reach an average ASR above $87\%$ across five recent safety-aligned models—exceeding existing baselines by over 15%. Under the black-box setting, DTA still outperforms its counterparts by over 25%.
2 Background
2.1 Paradigms in White-box Jailbreaking
Jailbreaking aims to craft prompts that bypass an LLM’s safety alignment, inducing it to generate harmful or restricted content. In the white-box setting, the foundational goal is to optimize an adversarial suffix $S$, appended to a user’s harmful prompt $P$, to maximize the conditional probability of the target LLM generating a predefined, affirmative target prefix (e.g., “Sure, here is …”) $T=(t_{1},\cdots,t_{k})$. This objective can normally be formulated as minimizing the negative log-likelihood (or cross-entropy loss):
$S^{*}=\arg\min_{S}\mathcal{L}(S)=\arg\min_{S}\left(-\sum_{i=1}^{k}\log p(t_{i}\mid P,S,t_{<i})\right).$ (1)
Existing methods, such as GCG [6] and COLD-Attack [7] are prominent implementations of this paradigm, employing token-level and global-level optimization processes, respectively, to find an effective $S^{*}$.
However, recent researchers have recognized that constraining the optimization to a single, fixed target prefix is overly restrictive and can lead to inefficient optimization. To mitigate this limitation, methods like AdvPrefix [8] are proposed. The core idea of AdvPrefix is to relax the objective by introducing a set of multiple, curated target prefixes, $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{m}\}$. The optimization goal then becomes maximizing the probability of generating any of these targets, often approximated by minimizing the loss of the most likely target in the set:
$S^{*}=\arg\min_{S}\left(\min_{T_{j}\in\mathcal{T}}\left(-\sum_{i=1}^{|T_{j}|}\log p(t_{j,i}\mid P,S,t_{j,<i})\right)\right)$ (2)
This multi-target approach provides more flexibility but still relies on a predefined set of external targets. Furthermore, researchers have proposed jailbreaking methods based on Reinforcement Learning (e.g., RLbreaker [9]) and adaptive jailbreaking methods (e.g., llm-adaptive-attacks [10]. However, all existing methods rely on externally-defined objectives, which may not adaptively align with different target LLMs’ natural harmful responses.
2.2 Language Model Decoding and Sampling Strategies
Given a prompt $x$, the Large Language Model (LLM) generates a response by autoregressively sampling tokens from a conditional probability distribution, i.e.,
$p_{T}(y^{i}_{t}\mid y_{<t},x)=\frac{e^{f^{i}(x,y_{<t})/T}}{\sum_{j}e^{f^{j}(x,y_{<t})/T}},$ (3)
where $y_{t}$ refers to the token output at position $t$ and $y_{<t}$ denotes the generated sub-sequence before $t$. In general, there are two commonly-used strategies for sampling $y_{t}$, i.e., greedy sampling and random sampling. For greedy sampling, the LLM selects the next token at position $t$ by $i=\arg\max_{i}p(y^{i}_{t}\mid y_{<t},x)$. For random sampling, the LLM samples the token according to the probability distribution in Eq. 3 with a non-zero temperature $T$.
In practice, most LLMs adopt greedy sampling or random sampling with a low temperature, which forms a narrow output space. A safety-aligned LLM usually generates refusal responses from this narrow space to harmful prompts. However, recent work [11] indicates that a broader output space randomly sampled from a safety-aligned LLM usually contains harmful responses. Inspired by this observation, we propose DTA to exploit the potentially harmful responses randomly sampled from the LLM as targets to optimize the adversarial suffix, aiming to induce similar responses under the modes of greedy sampling or random sampling with a small temperature.
3 Methodology
3.1 Overview of Dynamic Target Attack
The design of DTA begins from a simple intuition: optimizing the adversarial suffix $S$ toward harmful responses that have relatively high probability in the target model’s conditional output distribution is more efficient than using a fixed, relatively low probability response as the target. Given a harmful prompt $P$, DTA learns an adversarial suffix $S$ such that the target LLM produces a complete harmful response under standard decoding. As shown in Figure 2, DTA proceeds in short exploration–optimization cycles:
  1. 1.
    Dynamic Target Exploration. Temporarily relax decoding to explore a broader output region of the target model and directly sample native harmful responses; select the most harmful as the dynamic target $r^{*}$.
  2. 2.
    Target-Conditioned Optimization. Update $S$ with a small number of gradient optimization iterations to increase the likelihood of producing $r^{*}$ under standard decoding while preserving suffix fluency and avoiding refusal triggers.
Each cycle updates $r^{*}$ through re-sampling, which (i) anchors optimization in high-density regions of the current conditional distribution and (ii) solves the problem of distribution shift induced by optimizing $S$. Formally, the optimization in each cycle solves
$\min_{S}\ \mathcal{L}_{\text{DTA}}(P,S;r^{*})$ $=\mathcal{L}_{\text{resp}}(P,S;r^{*})+\lambda\cdot\mathcal{L}_{\text{suffix}}(S),$ (4)
where $\mathcal{L}_{\text{resp}}$ steers the model toward $r^{*}$ and $\mathcal{L}_{\text{suffix}}$ is a regularizer that promotes fluency while penalizing proximity to refusal-related tokens. The scalar $\lambda>0$ balances these terms.
Figure 2: Overview of DTA. DTA progressively executes an “sampling-optimization cycle” to directly sample the inherent harmful response from the target LLM’s relatively high-probability generation regions and optimize the adversarial suffix. Algorithm 1 shows the details of our DTA.
3.2 Exploration — Dynamically Sampling a Harmful Target
The cornerstone of DTA is that targets are directly sampled from the model’s conditional output distribution, rather than a fixed, low-probability response. Given a harmful prompt $P$ and the current suffix $S$, we query the target model $f$ with a relaxed, high-entropy decoding strategy that increases sampling diversity, thereby directly sampling candidate responses from the target model. Conditioned on P+SP{+}S, we draw $N$ candidates $\{{r_{i}}\}_{i=1}^{N}$ from the target model $f$:
$r_{i}\sim f\big(\cdot\mid P{+}S;\ \text{relaxed decoding}\big),\quad i=1,\ldots,N.$ (5)
Each $r_{i}$ is scored by a harmfulness judge $J$ (e.g., GPTFuzzer), yielding $J(r_{i})\in[0,1]$. We select the most harmful candidate as the cycle’s dynamic target:
$r^{*}=\arg\max_{r_{i}}J(r_{i}).$ (6)
By anchoring optimization to $r^{*}$—a high-density, model-native harmful response—DTA aligns its objective with what the target model already deems plausible, improving its efficiency and stability.
Algorithm 1 Dynamic Target Attack (DTA)
0:  Target LLM $f$, Reference LLM $f_{ref}$, Harmful prompt $P$, Exploration cycles $M$, Optimization iterations per cycle $T$, Samples per cycle NN, Learning rate $\eta$, Harmfulness threshold $\theta_{stop}$
1:  Initialize adversarial suffix $S$ randomly
2:  for $m=1$ to $M$ do
3:   /* — Phase 1: Sampling — */
4:   Sample NN candidate responses $\{r_{i}\}_{i=1}^{N}$ from $f_{ref}(P+S)$ using high temperature $\tau_{\text{search}}$.
5:   Evaluate harmfulness scores $\{J(r_{i})\}_{i=1}^{N}$ using a judge model.
6:   Select dynamic target response $r^{*}=\arg\max_{r_{i}}J(r_{i})$. {See Eq. 6}
7:   Truncate target to a fixed length for optimization: $r^{*}_{L}$.
8:   /* — Phase 2: Optimization — */
9:   for $t=1$ to $T$ do
10:    // Compute losses based on the target LLM $f$ with standard temperature $\tau_{\text{eval}}$
11:    Compute response loss $\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})$ over the truncated target. {See Eq. 7}
12:    Compute suffix regularization $\mathcal{L}_{\text{suffix}}(S)=\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S)$. {See Eqs. 8, 9}
13:    Compute total DTA loss $\mathcal{L}_{\text{DTA}}=\mathcal{L}_{\text{resp}}+\lambda\mathcal{L}_{\text{suffix}}$. {See Eq. 11}
14:    Update suffix via gradient descent: $S\leftarrow S-\eta\cdot\nabla_{S}\mathcal{L}_{\text{DTA}}(P,S)$. {See Eq. 12}
15:    // Check for early stopping condition
16:    Generate a test response $r_{test}=f(P+S)$ under standard decoding.
17:    if $J(r_{test})>\theta_{stop}$ then
18:     break {Terminate refinement early if successful}
19:    end if
20:   end for
21:  end for
22:  Return Optimized adversarial suffix $S^{*}$
3.3 Optimization — Efficiently Steering Toward the Dynamic Target
Given $r^{*}$, DTA updates $S$ via a small number of gradient optimization steps to minimize $\mathcal{L}_{\text{DTA}}$. This objective is composed of two primary components: a response likelihood loss $\mathcal{L}_{\text{resp}}$ and a suffix regularization term $\mathcal{L}_{\text{suffix}}$. Instead of matching the entire, potentially noisy sequence, we truncate $r^{*}$ to a fixed length of response, i.e., $r^{*}_{L}$. We employ this truncation strategy because (1) it mitigates the influence of noisy or incoherent tokens that often appear in the latter part of long generated responses, ensuring that the optimization focuses on more reliable and meaningful content, and (2) it reinforces early-stage control, encouraging the suffix to steer the model output from the very beginning of generation toward harmful behaviors.
The response-likelihood term is then the negative log-likelihood of $r^{*}_{L}$ under standard decoding (temperature $\tau_{\text{eval}}{=}0.7$):
$\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})=-\sum_{t=1}^{L}\log p_{\tau_{\text{eval}}}(y_{t}\mid y_{<t},,P{+}S).$ (7)
To preserve stealthiness and avoid triggering refusals, we add a suffix regularizer composed of the fluency loss $\mathcal{L}_{flu}$ and the refusal loss $\mathcal{L}_{rej}$:
$\mathcal{L}_{\text{flu}}(S)$ $=-\sum_{j=1}^{|S|}\log p_{\tau_{\text{eval}}}(s_{j}\mid s_{<j}),$ (8)
$\mathcal{L}_{\text{rej}}(S)$ $=\sum_{j=1}^{|S|}\ \sum_{v\in\mathcal{V}_{\text{rej}}}\log p(s_{j}{=}v\mid s{<j}),$ (9)
where $\mathcal{V}_{\text{rej}}$ is a set of common refusal tokens/phrases. We combine them as
$\mathcal{L}_{\text{suffix}}(S)=\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S),$ (10)
so that optimization jointly improves fluency while moving away from refusal-inducing regions in logit space. The overall objective is
$\mathcal{L}_{\text{DTA}}(P,S;r^{*})=\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})+\lambda\cdot(\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S)),$ (11)
and we update $S$ by gradient descent (Adam [12] in practice):
$S\leftarrow S-\eta\nabla{S}\mathcal{L}_{\text{DTA}}(P,S;r^{*}).$ (12)
3.4 Iterative Re-Sampling
We note that the first sampled target in a cycle can still be safe. In that case, continued optimization toward this safe target may stall progress and fail to jailbreak the target model. For example, in Figure 3, the first target $T^{1}_{S}$ is more harmful than the initial refusal response $R_{\text{initial}}$ but remains safe, and the first update still yields a safe output $R^{1}_{S}$. After each short optimization round, DTA re-samples from the model’s current conditional distribution given the harmful prompt and the updated suffix. It then selects the most harmful candidate as the next temporary target. This refresh keeps the target anchored to what the model now considers plausible, and it is repeated every cycle rather than focusing to a single target.
Figure 3: Illustration of the core difference between DTA and existing methods.
The brief update toward a model-native target shifts probability distribution away from refusal and reduces the target–distribution discrepancy, which increases the chance that the next re-sampling step surfaces a genuinely harmful, higher-density target. In Figure 3, the second re-sampled target $T^{2}_{S}$ is more harmful than $T^{1}_{S}$. In the second round, DTA optimizes the adversarial suffix $S$ towards $T^{2}_{S}$, successfully inducing the harmful response $R^{2}_{S}$. By alternating short updates with re-sampling, DTA stays aligned with the evolving output space, shortens the optimization path, and converges faster to an effective adversarial suffix. Empirically, this pattern matches Table 4, where increasing the number of re-sampling cycles improves ASR under a fixed compute budget.
4 Experiments
4.1 Experimental Setup
Target Models. Following prior work [7, 13, 8], we evaluate DTA against five popular open-source LLM families: Llama-3-8B-Instruct (Llama-3) [2], Vicuna-7B-v1.5 (Vicuna) [14], Qwen2.5-7B-Instruct (Qwen2.5) [3], Mistral-7B (Mistral) [15], and Gemma-7B (Gemma) [16]. These models represent diverse architectures and alignment strategies, enabling a comprehensive evaluation and comparison with existing methods. In our default white-box setting, the reference model ($f_{ref}$) is a copy of the target model ($f$). In our black-box setting, we employ two small models, i.e., Llama-3.2-1B-Instruct and Llama-3-8B-Instruct as the surrogate models, to jailbreak two large-scale LLMs, i.e., Llama-3-70B-Instruct and Kimi-K2-Instruct.
Datasets. We conduct experiments on two standard jailbreak benchmarks: AdvBench [6] and HarmBench [17]. To ensure consistent and efficient evaluation while maintaining diversity, we follow previous work [7, 6] and randomly sample a testing subset of 100 prompts from each dataset.
Baselines. Under the white-box setting, we compare DTA against six state-of-the-art white-box methods:COLD-Attack [7], GCG [6], llm-adaptive [10], I-GCG [13], AdvPrefix [8], and RLbreaker [9]. To ensure a fair and rigorous evaluation, all baselines are run with the hyperparameters recommended in their public implementations. Under the black-box setting, we compare DTA against three state-of-the-art black-box methods: ReNeLLM [18], PAP [19], and TAP [20]—again following their default configurations. In addition, we also select two representative white-box methods, COLD-Attack and AdvPrefix, to compare DTA’s performance with their transferability results.
Evaluation metrics. Following previous work [7, 13, 21], we employ a multi-faceted evaluation strategy using three distinct metrics to comprehensively measure attack performance: ${ASR}_{G}$ (judged by GPTFuzzer [22]), ${ASR}_{O}$ (judged by GPT-4), and ${ASR}_{L}$ (judged by Llama-Guard-3-8B [2]). Given a harmful prompt $P$ and the induced response $R$, a judge $\mathcal{J}$ returns an unsafe score. For GPTFuzzer and Llama-Guard-3-8B the score lies in $[0,1]$, denoted as $\mathcal{J}(P,r)$, and we compute $\text{ASR}_{G}$ and $\text{ASR}_{L}$ as
$\text{ASR}_{G}\text{ or }\text{ASR}_{L}=\frac{\sum_{(P,R)}\mathbb{I}(\mathcal{J}(P,R)\geq 0.5)}{\#\text{Total Count}},$ (13)
where $\mathbb{I}$ is a indicator function and we set the threshold as $0.5$.
When using GPT-4 as the judge, we pass a standardized judge template and the response $R$ to obtain a discrete unsafe score $\mathcal{J}(\text{template},R)\in[1,5]$. We report $\text{ASR}_{O}$ by thresholding this score at $4$:
$\text{ASR}_{O}=\frac{\sum_{R}\mathbb{I}(\mathcal{J}(\text{template},R)\geq 4)}{\#\text{Total Count}}.$ (14)
We provide further details in the Appendix.
Implementation Details. In each sampling phase, we sample $N=30$ candidate responses from the target LLM. In the optimization phase, we set the reference model under a standard decoding strategy, e.g., a low decoding temperature of $\tau_{eval}=0.7$ to ensure stable generation. For each of the $M=20$ sampling rounds, the adversarial suffix is optimized for $T=10$ iterations using the Adam optimizer. We apply an early stopping condition, terminating the process for a given prompt if a generated response achieves a harmfulness score above 0.9. All experiments are conducted on two NVIDIA RTX A6000 GPUs, each with 48 GB of memory.
4.2 Main Results under the White-box Setting
Table 1: Comparison of attack success rates (ASRs) achieved by DTA and baselines on five open-sourced LLMs across two datasets. We set M = 20, T = 10 for DTA. We follow the recommended settings for six baselines to achieve their best performance. The red values indicate the highest ASRs.
Method Llama-3 Vicuna Qwen2.5 Mistral Gemma
$\text{ASR}_{G}$ $\text{ASR}_{O}$ $\text{ASR}_{L}$ $\text{ASR}_{G}$ $\text{ASR}_{O}$ $\text{ASR}_{L}$ $\text{ASR}_{G}$ $\text{ASR}_{O}$ $\text{ASR}_{L}$ $\text{ASR}_{G}$ $\text{ASR}_{O}$ $\text{ASR}_{L}$ $\text{ASR}_{G}$ $\text{ASR}_{O}$ $\text{ASR}_{L}$
AdvBench
COLD-Attack 43% 64% 56% 33% 51% 41% 23% 26% 22% 73% 81% 69% 49% 58% 55%
GCG 25% 47% 56% 15% 28% 22% 19% 37% 45% 46% 77% 67% 19% 37% 34%
llm-advaptive 92% 97% 90% 72% 91% 85% 92% 91% 87% 90% 95% 92% 95% 97% 89%
I-GCG 20% 85% 94% 47% 97% 93% 15% 56% 52% 20% 89% 86% 12% 29% 46%
AdvPrefix 75% 93% 87% 80% 93% 88% 79% 90% 88% 86% 93% 86% 72% 78% 83%
RLbreaker  4%  6% 16% 57% 76% 70% 49% 50% 54% 81% 94% 89% 79% 90% 87%
DTA 92% 98% 94% 94% 93% 94% 81% 59% 80% 95% 97% 93% 92% 99% 75%
HarmBench
COLD-Attack 28% 42% 46% 18% 29% 37% 23% 37% 31% 42% 58% 50% 38% 46% 44%
GCG 14% 37% 56%  9% 23% 25% 17% 40% 51% 29% 55% 57%  6% 17% 21%
llm-advaptive 62% 68% 91% 50% 66% 91% 59% 70% 90% 52% 66% 91% 67% 69% 76%
I-GCG 17% 73% 96% 29% 73% 86% 20% 48% 63% 19% 75% 89% 13% 19% 37%
AdvPrefix 44% 67% 88% 43% 73% 75% 44% 63% 79% 48% 63% 76% 46% 53% 71%
RLbreaker  6% 16% 29% 53% 76% 78% 34% 47% 59% 52% 71% 78% 45% 65% 59%
DTA 68% 85% 96% 64% 80% 96% 65% 75% 97% 65% 84% 97% 62% 88% 44%
Figure 4: Comparison result of DTA and baselines on HarmBench. Dark (Light) bar denotes the average (maximum) ASRs across five target LLMs.
We conduct extensive white-box attack experiments on two widely used benchmarks, AdvBench and HarmBench, across five recent safety-aligned LLMs. In Table 1, we report the detailed ASR results under three evaluation metrics (i.e., GPTFuzzer, GPT-4, and Llama-Guard-3-8B). Figures 5 and 4 show the average (dark bars) and maximum (light bars) ASRs of DTA and existing methods to highlight both effectiveness and stability. Overall, the results clearly show that DTA consistently outperforms existing baselines across two datasets in most cases.
[…]
Figure 5: Comparison result of DTA and baselines on AdvBench. Dark (Light) bar denotes the average (maximum) ASRs across five target LLMs.
Beyond the outperforming performance of ASRs, DDTA also exhibits stable behavior across the five target LLM families. On HarmBench, for instance, I-GCG could achieve above 70% $\text{ASR}_{O}$s on Llama-3, Vicuna and Mistral, on Llama-3, Vicuna, and Mistral, yet drops to just 19% on Gemma. A plausible reason is that I-GCG optimizes toward a long, predefined response that typically resides in an extremely low-density region of the model’s conditional output space, which amplifies the target–distribution discrepancy and makes outcomes highly model-dependent. In contrast, under the white-box setting, DTA delivers consistently high performance across diverse LLMs by anchoring its targets in relatively high-density regions via dynamic sampling and iterative re-sampling.
4.3 Results under the Black-box Setting
Under the black-box setting, i.e., only response access to the target model is available, DTA also exhibits its effectiveness in jailbreaking two large LLMs (Llama-3-70B-Instruct and Kimi-K2-Instruct). As shown in Table 2, using Llama-3-8B-Instruct as a surrogate model to jailbreak Llama-3-70B-Instruct, DTA achieves an ASR of 85%, exceeding COLD-Attack (26%) and AdvPrefix (59%) by over 25%. Three black-box jailbreaking methods, ReNeLLM, PAP, and TAP achieve an ASR of 7%, 11%, and 11%, respectively. Against Kimi-K2-Instruct under the same local configuration, DTA reaches 70% ASR, which is competitive with AdvPrefix (72%) and higher than COLD-Attack and three black-box methods. Even with a much smaller local model (Llama-3.2-1B-Instruct), DTA achieves an ASR 30% against Llama-3-70B-Instruct and an ASR of 50% against Kimi-K2-Instruct, respectively. Overall, these results show that DTA remains highly effective in a black-box setting compared to both white-box and black-box methods.
Table 2: Transferability results of DTA across different local LLMs and target (reference) LLMs. In this table, the experiments are conducted on AdvBench and the Attack Success Rates (ASRs) are evaluated by GPTFuzzer (i.e., $\text{ASR}_{G}$). For white-box methods, we employ a local LLM as a surrogate for optimization.
Dataset        Local LLM $\rightarrow$ Target LLM White-box Black-box DTA
COLD-Attack AdvPrefix ReNeLLM PAP TAP
AdvBench
Llama-3.2-1B-it $\rightarrow$ Llama-3-70B-it   7% 53%   7% 11% 11% 30%
Llama-3-8B-it $\rightarrow$ Llama-3-70B-it 26% 59% 85%
Llama-3.2-1B-it $\rightarrow$ Kimi-K2-it   4% 72% 50% 49% 36% 50%
Llama-3-8B-it $\rightarrow$ Kimi-K2-it 20% 72% 70%
4.4 Response diversity
Figure 6: Examples of generated responses by DTA and target responses. As exemplified by these examples, DTA could select diverse responses as targets, which is more harmful than the predefined response. DTA could generate harmful and clear responses.
As shown in Figure 6, DTA dynamically explores and discovers harmful responses to serve as its targets. These targets are fundamentally superior to the static, predefined prefixes (e.g., “Sure, here is…”) used by existing methods, as they are not generic affirmations but diverse, context-specific responses already aligned with the harmful goal. These targets range from conceptual explanations of psychological tactics to structured, step-by-step guides for creating phishing emails or selling illegal drugs. Crucially, these targets are not merely aligned with the prompt; they are often substantially more harmful in themselves than a simple affirmative phrase. By optimizing towards these high-quality, empirically-grounded targets, DTA efficiently refines the adversarial suffix and successfully jailbreaks the target LLMs.
4.5 Efficiency and Iteration Cost
Table 3: Iteration and Running time costs. We follow the existing work’s default settings to evaluate their performance. RLbeaker trains an agent and has no optimization iterations.
Method Iteration Time Cost (h)
COLD-Attack 2,000 10.2
GCG 1,000 53.6
I-GCG 500 24.5
llm-adaptive 10,000 8.5
AdvPrefix 1,000 23.7
RLbreaker 10.0
DTA 200 3.7
In Table 3, we present a comparison of the iteration and time costs for DTA and several baselines. The results reveal that existing methods, whether based on fixed-target optimization or more advanced adaptive strategies, often require a substantial number of iterations, typically ranging from 1,000 for methods like GCG and AdvPrefix, to as many as 10,000 for certain adaptive attacks like llm-adaptive-attacks. This high cost stems from the inherent difficulty of forcing a model towards an external, and often unnatural, target distribution.
[…]
4.6 Hyper-parameter selection
We conduct an experimental study to explore the impact of the exploration round $M$ and the per-round optimization steps $T$ on DTA’s performance. In Table 4, we report the experimental results of for jailbreaking Llama-3-8B-Instruct on AdvBench while varying the values of $M$ and $T$ under a roughly fixed per-prompt budget $M\times T=200$). The experimental results show that allocating more computational budget to exploration consistently and dramatically improves the ASR.
Table 4: Experimental Study on the impact of dynamic sampling. The light green column represents DTA’s default settings (i.e., M = 20, T = 10). We set the target and reference LLM as Llama-3-8B-Instruct. We employ GPTFuzzer to evaluate Attack Success Rates (i.e., $\text{ASR}_{G}$).
Attack Success Rates on AdvBench (M = Sampling Iterations, T = Optimization Iterations)
M = 1 M = 2 M = 4 M = 5 M = 10 M = 20 M = 20 M = 40 M = 100 M = 200
T = 200 T = 100 T = 50 M = 40 T = 20 T = 5 T = 10 T = 5 T = 2 T = 1
34% 41% 65% 70% 85% 89% 92% 95% 97% 100%
Specifically, under the optimization-heavy setting, i.e., sampling once ($M=1$) and optimizing 200 steps ($T=200$, DTA only achieves an ASR of 34%. However, when we increase $M$ such as $M=10$ and $T=20$, DTA significantly improves its performance, achieving an ASR of 85%. Moreover, if we consider an extreme situation, i.e., allocating all budget to exploration ($M=200$) and only optimizing one step ($T=1$) each round, DTA could achieve an ASR of 100%. These results yield an insight: when the chosen target response lies far from high-density regions of the model’s conditional output distribution, additional gradient steps yield little improvement. DTA allocates more of the fixed budget to exploration to resample targets nearer to these high-density modes, thereby explicitly reducing this discrepancy. Once the optimization objective is aligned with the model’s native output landscape, each gradient step becomes more productive, the target likelihood rises faster, and the same compute delivers substantially higher ASR.
5 Conclusion
In this paper, we propose Dynamic Target Attack (DTA), a novel jailbreak method that relies on directly sampling candidate responses as targets to optimize the adversarial prompt. In each optimization round, DTA iteratively samples multiple responses and selects the most harmful one as the current round target. In contrast to existing methods, DTA selects targets from the target model’s relatively high-density region, significantly reducing the discrepancy between the target and the model’s output distribution, thereby accelerating convergence to an effective adversarial prompt. Throughout our extensive experiments on diverse jailbreak benchmarks, DTA demonstrates its effectiveness against multiple recent safety-aligned LLMs under both white- and black-box settings, consistently outperforming the state-of-the-art baselines.
Appendix A Related Work
Here, we introduce six white-box and three black-box methods as our baselines in this paper:
A.1 White-box baselines
GCG [6] (Greedy Coordinate Gradient) is a foundational white-box attack. It employs a greedy, coordinate descent-based optimization to find an adversarial suffix $S$ that, when appended to a harmful prompt $P$, maximizes the probability of the model generating a fixed, predefined target response (e.g., “Sure, here is…”).
COLD-Attack. To address the limitations of high-perplexity attacks like GCG, Guo et al. (2024) introduced COLD-Attack, a framework that formally reframes controllable jailbreaking as a controllable text generation problem. By adapting the COLD algorithm, this method operates in the continuous logit space rather than on discrete tokens. It defines a composite energy function to unify multiple differentiable constraints, such as attack success, fluency, and semantic similarity. The framework uses Langevin dynamics to sample optimal logits that balance these competing objectives, and then employs a specialized LLM-guided decoding process to ensure the final text is coherent. This unified approach enables the generation of highly stealthy, low-perplexity attacks and introduces novel, controllable scenarios like paraphrasing and position-constrained attacks, which are not achievable by previous methods.
llm-adaptive [10] is an adaptive methodology for efficiently generating successful jailbreaking attacks against state-of-the-art, safety-aligned large language models. Operating under gray-box or black-box assumptions, its core strategy combines customized, handcrafted prompt templates with simple optimization algorithms, like random search, to iteratively refine attack strings based on model feedback. A key component of its effectiveness is the "self-transfer" technique, where successful attack patterns discovered on one harmful behavior are used to rapidly accelerate the attack on others, and it further exploits model-specific vulnerabilities, such as optimizing against token log-probabilities or leveraging unique API features like prefilling, to bypass safety filters with surprisingly few queries.
I-GCG. Jia et al. [13] proposed I-GCG, enhancing optimization-based jailbreaks by addressing the limitations of GCG. They found that GCG’s simple target template (e.g., "Sure, here is…") was insufficient and instead introduced diverse target templates containing harmful self-suggestions (e.g., "my output is harmful") to more effectively mislead the LLM. To improve optimization efficiency, they developed an automatic multi-coordinate updating strategy that adaptively replaces multiple tokens per step, accelerating convergence. Additionally, they employed an easy-to-hard initialization technique, using a suffix generated from a simple malicious query as the starting point for more difficult ones.
AdvPrefix [8] is an optimization objective used in gradient-based adversarial attacks like GCG to "jailbreak" Large Language Models (LLMs). The goal is to find a short, often nonsensical, string of characters (the "prefix") that, when attached to a user’s prompt, hijacks the model’s output generation. This objective function is optimized to find the prefix that maximizes the probability of the LLM generating a specific, malicious target response (e.g., harmful instructions or hate speech), thereby bypassing its safety alignments.
RLbreaker [9] is an automated jailbreak prompt generation method that leverages Deep Reinforcement Learning (DRL) to systematically bypass the safety alignments of Large Language Models (LLMs). The system features three components: a DRL agent (attacker) that iteratively generates and refines prompts, a target LLM (victim), and an evaluator LLM (judge) that provides a reward signal based on the attack’s success. During training, the agent continuously modifies a prompt (up to 5 steps) and attacks the target, learning from the evaluator’s feedback through this trial-and-error process to master strategies for creating effective jailbreak prompts.
A.2 Black-box baselines
ReNeLLM. To address the limitations of manual and optimization-based jailbreaks, Ding et al. [18] proposed ReNeLLM, an automated framework that leverages LLMs themselves to generate effective and transferable jailbreak prompts. Their method generalizes the attack into two main stages: (1) Prompt Rewriting, which automatically applies a series of linguistic transformations (such as paraphrasing, misspelling sensitive words, or partial translation) to disguise the harmful intent of an initial query, and (2) Scenario Nesting, which embeds the rewritten prompt into a seemingly benign and common instruction-following scenario, such as code completion, table filling, or text continuation. This combined approach is shown to effectively shift the target LLM’s execution priority from safety to usefulness, enabling it to bypass safeguards and achieve high attack success rates on state-of-the-art models like GPT-4 and Claude-2.
PAP. Zeng et al. [19] introduce a novel jailbreak methodology, Persuasive Adversarial Prompts (PAPs), which reframes the attack vector by treating LLMs as human-like communicators susceptible to social influence rather than as mere instruction followers. Their approach is systematically grounded in a comprehensive taxonomy of 40 persuasion techniques derived from decades of social science research. The method employs a two-stage framework: first, they train a "Persuasive Paraphraser" by fine-tuning an LLM (e.g., GPT-3.5); second, this trained model is deployed to automatically generate new, human-readable PAPs at scale. The effectiveness of these prompts in bypassing safety alignments is then systematically evaluated across various LLMs, including Llama 2 and the GPT series, using a GPT-4 Judge to classify the harmfulness of the outputs.
TAP. Mehrotra et al. [20] propose Tree of Attacks with Pruning (TAP), an automated, black-box method for generating interpretable jailbreaks for LLMs. TAP utilizes an attacker LLM to iteratively generate and refine candidate prompts in a tree structure (branching) and an evaluator LLM to assess them. This process involves two pruning phases: first, pruning off-topic prompts before querying the target, and second, pruning low-scoring prompts after assessing the target’s response.
Appendix B Supplementary Materials on Experimental Setups
B.1 Target LLMs.
Here we briefly introduce the information of five open-sourced target LLMs:
  • Llama-3-8B-Instruct [2]: LLaMA-3-8B-Instruct is an instruction-tuned version of Meta’s LLaMA-3 language model with 8 billion parameters. It is trained to follow natural language instructions across a wide range of tasks and serves as a strong open-source alternative to proprietary chat models.
  • Vicuna-7B-v1.5 [14]: Vicuna-7B-v1.5 is a fine-tuned version of LLaMA-2-7B, developed by LMSYS, optimized for multi-turn conversational ability. It is trained on user-shared ChatGPT conversations and aims to match the quality of closed-source models such as GPT-3.5.
  • Qwen-2.5-7B-Instruct [3]: Qwen-2-7B-Instruct is the instruction-tuned variant of Alibaba’s Qwen-2-7B model, trained to follow prompts and generate aligned responses. It supports multilingual understanding and demonstrates competitive performance on safety and alignment benchmarks.
  • Mistral-7B-Instruct-v0.3 [15]: Mistral-7B-Instruct is an instruction-tuned version of the Mistral-7B model released by Mistral AI. It features efficient inference via grouped-query attention and sliding-window attention, and has been tuned on a wide range of instruction-following tasks.
  • Gemma-7b [16]: Gemma is Google’s family of lightweight, state-of-the-art open models, built on the same research and technology as Gemini.
Table 5 shows the source links of five target LLMs we used in this paper.
Table 5: Source repository of five target LLMs.
Model Name Link
Llama-3-8B-Instruct https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
Vicuna-7B-v1.5 https://huggingface.co/lmsys/vicuna-7b-v1.5
Qwen-2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Mistral-7B-Instruct-v0.3 https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3
Gemma-7b https://huggingface.co/google/gemma-7b
B.2 Judge Models
Table 6: Source repository of two open-sourced judge models.
Model Name Link
GPTFuzzer https://huggingface.co/hubert233/GPTFuzz
Llama-Guard-3-8B https://huggingface.co/meta-llama/Llama-Guard-3-8B
Table 6 shows the links for downloading GPTFuzzer and Llama-Guard-3-8B. Below, we briefly introduce two judge models:
  • GPTFuzzer [22] is a fine-tuned RoBERTa-based model designed to evaluate the harmfulness of generated text. It is trained to classify responses as either “safe” or “unsafe”, providing a probability score that indicates the likelihood of the input text being harmful. As the results reported in [22]’s Table 1, GPTFuzzer achieves superior performance compared to traditional methods, offering both high efficiency and effectiveness in detecting potentially harmful content in LLMs.
  • Llama-Guard-3-8B [2] is a fine-tuned version of Meta’s Llama-3-8B model, specifically designed to serve as a protective guardrail model to prevent the generation of harmful or unethical content by identifying and blocking such outputs. It supports multilingual harmful content detection and classifies content into 14 distinct harmful categories, providing robust protection for developers by identifying various types of illegal or unethical content.
B.3 Benchmark Datasets
In this paper, we employ two widely-used benchmark datasets to comprehensively evaluate the effectiveness of DTA compared to baselines. Here we briefly introduce two datasets and Table 7
  • AdvBench is a widely adopted benchmark dataset designed for evaluating adversarial robustness in LLMs. It includes a variety of harmful queries, such as requests for instructions on illegal activities, and is used to test a model’s ability to resist malicious prompts. AdvBench evaluates how well models can handle adversarial inputs without generating harmful content.
  • HarmBench is another comprehensive dataset that focuses on harmful instructions and unethical requests. It includes prompts that involve a wide range of harmful scenarios, such as hacking, abuse, and other illegal or morally questionable activities. HarmBench is intended to evaluate how models handle potentially harmful outputs and whether they can resist generating content that violates ethical guidelines.
Table 7: Source repository of two benchmark datasets.
Dataset Link
AdvBench https://github.com/llm-attacks/llm-attacks
HarmBench https://huggingface.co/datasets/JailbreakBench/JBB-Behaviors
B.4 Judge Template for GPT-4
In this paper, we employ GPT-4 as a judge model to evaluate the performance of DTA and existing baselines. Specifically, following previous work [13, 8], we input a judge template and a harmful prompt into GPT-4, generating an unsafe score. Here we provide the judge template:
Figure 7: Judge template for GPT-4 as a judge.
Appendix C Supplementary Experimental Settings
C.1 Response Generation Settings
We set the temperature to 0.7 following previous work [7], the maximum number of token sequences to 256, and top_k to 50. We show the generation pseudo-code as follows:
1def response_generation_from_huggingface_llms(
2 llm: transformers.AutoModelForCausalLM, tokenizer: transformers.AutoTokenizer, prompt: str, temperature: float = 0.7, top_p: float = 0.95, top_k: int = 50, max_n_tokens: int = 256, num_return_sequences: int = 1,
3):
4 inputs = tokenizer(prompt, return_tensors = "pt").to(llm.device)
5 responses = llm.generate(
6 **inputs,
7 temperature = temperature,
8 top_p = top_p,
9 top_k = top_k,
10 num_return_sequences=num_return_sequences,
11 max_new_tokens = max_n_tokens,
12 )
13 responses = tokenizer.batch_decode(
14 outputs[:, input_len:],
15 skip_special_tokens = True,
16 )
17 return responses
18
19def response_generation_from_api(
20 model_name:str, prompt: str, temperature: float = 0.7, top_p: float = 0.95, top_k: int = 50, max_n_tokens: int = 256, num_return_sequences: int = 1,
21):
22 client = OpenAI(api_key = os.getenv("OPENAI_API_KEY"))
23 response = client.chat.completion.create(
24 model = model_name, messages = [{"role":"user", "content":prompt}], max_tokens=max_n_tokens, temperature=temperature, top_p=top_p, top_k=top_k, n = num_return_sequences,
25 )
26 response = [choice.message.content for choice in response.choices]
27 return response
Appendix D Open Source
Due to the purpose of academic research and improving the safety of LLMs, we open-source our code and necessary instructions to re-implement our experiment. We provide an anonymous repository at https://anonymous.4open.science/r/Dynamic-Target-Attack-4176.